# The Price of Uncertainty for Social Consensus

### v2

@@ 4.1 Proof of Improved Upper Bound
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon n^{2}\right).$
Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$ because we must begin with that many bad edges. This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
@@
$\Theta\left(m|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}m^{2}\right)$.
@@
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(n^{3/2}|S_{0}|\right).$
@@
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}n^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(n^{-1/4}\right)^{-2}n^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$
The result of Lemma 16 can be plugged into Eq. 7 with Proposition 11 to obtain a bound on $\sum E_{m}$. Since $\varepsilon=\Omega(n^{-1/4})$, we have
$\sum A_{m}$ $=\sum A_{0}+\varepsilon\sum_{i=1}^{m}z_{i}$
$\sum A_{m}$ $=O\left(n|S_{0}|\right)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$

### v3

@@ 4.1 Proof of Improved Upper Bound
$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\Theta(\varepsilon n)}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon n^{2}\right).$
Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$ because we must begin with that many bad edges. This means that $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ must include $O\left(\varepsilon n^{2}\right)$, and we are done with this case.
@@
$\Theta\left(n|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}n^{2}\right)$.
@@
$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}z_{j}$ $\leq\frac{\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}z_{j}$ $=O\left(n^{3/2}|S_{0}|\right).$
@@
$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}n^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon^{2}\left(n^{-1/4}\right)^{-2}n^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}z_{j}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right).$
The result of Lemma 16 can be plugged into Eq. 7 with Lemma 7 and Proposition 11 to obtain a bound on $\sum A_{m}$. Since $\varepsilon=\Omega(n^{-1/4})$, we have
$\sum A_{m}$ $=\sum A_{0}+m+\varepsilon\sum_{i=1}^{m}z_{i}$
$\sum A_{m}$ $=O\left(n|S_{0}|\right)+\Theta(n)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$